# Learning to Segment for Vehicle Routing Problems

### v1

@@ Abstract
Iterative search heuristics are widely recognized as state-of-the-art for solving Vehicle Routing Problems (VRPs). In this work, we identify and exploit a critical observation: within these solvers, a large portion of the solution remains stable, i.e., unchanged across search iterations, causing redundant computations, especially for large-scale VRPs with long subtours. To address this, we pioneer the formal study of the First-Segment-Then-Aggregate (FSTA) decomposition technique to accelerate iterative solvers. Specifically, FSTA preserves stable solution segments during the search, aggregates nodes within each segment into fixed hypernodes, and focuses the search only on unstable portions. Yet, a key challenge lies in identifying which segments should be aggregated by FSTA. To this end, we then introduce Learning-to-Segment (L2Seg), a novel neural framework to intelligently differentiate potentially stable and unstable portions for FSTA decomposition. We present three L2Seg variants: non-autoregressive (globally comprehensive but locally indiscriminate), autoregressive (locally refined but globally deficient), and their synergy, with bespoke training and inference strategies. Empirical results on CVRP and VRPTW suggest that L2Seg accelerates state-of-the-art iterative solvers by up to 7x. Additionally, we provide in-depth analysis showing NAR and AR synergy achieves best performance by combining their complementary strengths. Notably, L2Seg is a flexible framework that is compatible with traditional, learning-based, and hybrid solvers, while supporting a broad class of VRPs.
@@ 1 Introduction
Vehicle Routing Problems (VRPs) have profound applications such as in logistics and ride-hailing, driving decades of advances in combinatorial optimization [1]. As NP-hard problems, they are typically tackled with heuristics approximately. Neural Combinatorial Optimization (NCO) [2, 3, 4, 5, 6] has recently introduced machine learning into VRP solving, enabling data-driven decision-making with minimal domain knowledge while matching and even surpassing the performance of meticulously designed heuristics such as Lin-Kernighan-Helsgaun (LKH) [7] and Hybrid Genetic Search (HGS) [8].
Generally, state-of-the-art classic and neural VRP solvers predominantly rely on iterative search to refine the current best-so-far solution through local search (e.g., ruin and repair). However, as we observe in Section 3, a significant portion of edges stabilize, or their presence in the solution stops changing between iterations, as the search progresses, despite repeated local search computations. Specifically, we refer stable edges as those that consistently remain in the solution across iterations, while unstable edges are likely to be re-optimized. This is particularly evident in VRPs with long subtours, where inner edges remain fixed while only boundary edges undergo frequent combinatorial changes. Intuitively, such stability can be inferred from the spatial distribution of customer nodes and the solution properties through end-to-end machine learning. Yet, existing solvers overlook such opportunities, leading to redundant computations that hinder their scalability and efficiency, especially in large-scale VRPs with long subtours.
@@
Yet, identifying unstable portions for effective FSTA decomposition remains a key challenge. We then introduce Learning-to-Segment (L2Seg), a novel learning-guided framework that leverages deep models to intelligently differentiate potentially stable and unstable portions, allowing dynamic decomposition for accelerated local search. Realizing this, however, is nontrivial: it involves a large combinatorial decision space, requiring accurate segment grouping across optimization stages. Moreover, it demands modeling complex interdependencies among predicted edges, constraints, spatial distribution, solution structures, and both node and edge features.
@@
To address these challenges, L2Seg considers a spatial locality strategy to restrict the search space and employs encoder-decoder-styled neural models. The encoder integrates graph-level and route-level features using attention and customized graph neural networks, generating node embeddings that guide edge re-optimization predictions in the decoder. L2Seg offers three decoder variants: (1) L2Seg-NAR (Non-Autoregressive): which features one-shot fast global prediction; (2) L2Seg-AR (Autoregressive), which enjoys sequential dependency modeling for high-precision local predictions; and (3) L2Seg-SYN (Synergized), which balances the strengths of NAR and AR for best performance. Notably, this represents a pioneering work that explores the joint decision-making between AR and NAR models in neural combinatorial optimization. Our L2Seg models are trained on datasets labeled using a lookahead procedure: edge stability is classified based on whether its presence in the solution was changed during re-optimization across the iterative search.
@@
Through extensive experiments on large-scale CVRPs and VRPTWs, we show that L2Seg accelerates heuristics by 2x to 7x, enabling them to outperform various state-of-the-art traditional, neural, and hybrid baselines. Notably, L2Seg exhibits strong flexibility and generality in enhancing a wide range of backbone solvers, including traditional LKH-3 [7] solver, other orthogonal Large Neighborhood Search (LNS) framework [9], and recent learning-guided decomposition method Learning-to-Delegate (L2D) [4]. We further analyze the synergy between AR and NAR models in segment decision-making. Our analysis reveals the complementary strengths and weaknesses of NAR and AR models, and shows their synergy achieves best performance by integrating NAR’s global detection capability with AR’s local refinement precision.
@@
Our contributions are as follows: (1) We make a critical yet underexplored insight that stable segments persist across search iterations in large-scale VRPs, causing redundant computations; (2) We formally study First-Segment-Then-Aggregate (FSTA), a generic decomposition framework that aggregates stable segments into fixed hypernodes while ensuring provable equivalence; (3) We propose Learning-to-Segment (L2Seg), a learning-guided framework with bespoke network architecture, training, and inference for identifying such segments; (4) We propose autoregressive, non-autoregressive, and their synergistic deep models, pioneering the first-of-its-kind study in NCO; (5) L2Seg consistently accelerates state-of-the-art iterative VRP solvers by 2x to 7x, boosting both traditional and learning-based solvers, including other decomposition frameworks. Our work presents a new perspective in NCO by introducing a segment-based acceleration framework that removes redundant search, achieves state-of-the-art on CVRP and VRPTW, while supporting a broad class of VRPs.
@@ 2 Related Works
Existing VRP Solvers. Classical VRP solvers include exact methods with guarantees [10] and practical heuristics [7]. Recently, inspired by Pointer Networks [11], machine learning has been applied to combinatorial optimization, either end-to-end [12, 2, 13, 14, 15, 16, 17, 18, 19] or following a learning-guided paradigm to inject data-driven insights into backbone optimizers [4, 20, 21, 22, 23, 24]. For VRPs, the former could yield competitive performance to classic methods [25, 16], while the latter often achieve state-of-the-art performance [4, 26]. Among these, most effective VRP solvers rely on iterative search, including classic heuristics such as HGS [8], LNS [9] and LKH [7]; neural solvers that learn local search [27, 28, 29, 30, 31]; neural constructive solvers integrated with search components [32, 25, 33, 34, 35, 36, 37]; and hybrid learning-guided methods like L2D [4]. However, both handcrafted and neural iterative search solvers overlook the redundant computations identified in this paper, particularly in large-scale VRPs.
@@
Decomposition for Large-scale VRPs. Scalability in VRP solvers often relies on effective decomposition [38]. Recently, machine learning has been leveraged to assist decomposition, such as sub-tour grouping [39, 4] and problem variant reduction [40, 41], or to boost neural solvers such as action space decomposition [16, 25, 42] and spatial-based decomposition [26, 43, 44]. In this paper, we present FSTA and L2seg, a fresh segment-based decomposition framework. Notably, L2Seg holds potential to enhance the other decomposition methods, such as LNS [9] and L2D [4]. While FSTA is conceptually related to path decomposition [38], which clusters nearby routes using barycenters, FSTA generalizes it by allowing segments of arbitrary length, reversals, and attribute aggregation. While another related work [45] explores segment stability for re-optimization in a specific dynamic CVRP setting, our work addresses a different problem, i.e., identifying stable segments across search steps to accelerate iterative solvers. And we formally analyze the nontrivial solution equivalence of FSTA across broader VRP variants. Moreover, L2Seg uniquely integrates three novel deep learning models (AR, NAR, and synergized) to dynamically guide FSTA decomposition during search.
@@
Autoregressive and Non-autoregressive Models. In NCO, NAR models make fast, global predictions such as edge selection heatmaps [34, 40, 46] in a one-shot manner. However, they struggle to model complex interdependencies, particularly VRP constraints. In contrast, AR models make sequential predictions, one variable at a time, offering stronger modeling capacity but might overlook global structure. They excel in complex tasks, e.g., learning-to-construct (e.g., [25]) and learning-to-search solvers (e.g., [28]). Recent NCO works combine AR and NAR models in divide-and-conquer frameworks, with NAR for sub-problem splitting and AR for solving [26, 40, 41]. Other than task separation, we are the first to leverage their complementary strengths for joint decision-making, enabling more effective identification of unstable segments in FSTA decomposition.
@@ 3.1 Vehicle Routing Problems (VRPs)
VRPs aim to minimize total travel costs (often distance or travel time) while serving a set of customers under constraints. Formally, A VRP $P$ is defined on a graph $G=\{V,E\}$, where each node $x_{i}\in V$ represents a customer and each edge $e_{i,j}\in E$ represents traveling from $x_{i}$ to $x_{j}$ and is associated with a travel cost. For Capacitated VRP (CVRP), vehicles of capacity $C$ start and end at a depot node $x_{0}$. The sum of the demands $d_{i}$ on any route must not exceed $C$, and each customer should be served exactly once. For VRP with Time Windows (VRPTW), each customer is additionally associated with a service time $s_{i}$ and a time window $\bigl{[}t^{l}_{i},\,t^{r}_{i}\bigr{]}$ within which service must begin.
@@ 3.2 FSTA Decomposition
Figure 1: Percentage of re-optimized edges during iterative search process: using LKH-3 on 100 CVRP instances of sizes 2k and 3k with large and small $C$. A vast majority of edges remain unchanged, suggesting redundant calculations.
@@
Figure 1 depicts that iterative search solvers perform redundant searches, reoptimizing only a small portion while many edges remain unchanged, especially in large subtours with high capacity $C$. More visualization are provided in Appendix A.1. Inspired by [45], we formally study the decomposition technique, First-Segment-Then-Aggregate (FSTA), for accelerating iterative search solvers. As shown in the top of Figure 2, FSTA segments the VRP solutions by identifying unstable portions, and then groups them into hypernodes with aggregated attributes. We thus expect more efficient re-optimization on the reduced problems.
@@
Segment Definition. Denote the solution (set of routes) of a CVRP as =$\mathcal{R}=\{R^{1},R^{2},...\}$, and each route as $R^{i}=(x_{0}\rightarrow x^{i}_{1}\rightarrow x^{i}_{2}\rightarrow...\rightarrow x_{0})\in\mathcal{R}$, where the first and the last nodes in $R^{i}$ are the depot. A segment consists of some consecutive nodes within a route. We denote the segment containing the $j^{\text{th}}$ to $k^{\text{th}}$ nodes of route $i$ as $S^{i}_{j,\,k}=(x_{j}^{i}\rightarrow...\rightarrow x_{k}^{i})$. An aggregated segment $\tilde{S}^{i}_{j,\,k}$ uses one hypernode ($\tilde{S}^{i}_{j,\,k}=\{\tilde{x}^{i}_{j,k}\}$) or two hypernodes ($\tilde{S}^{i}_{j,\,k}=\{\tilde{x}^{i}_{j},\tilde{x}^{i}_{k}\}$) with aggregated attributes (e.g. the demand of $\tilde{x}^{i}_{j,k}$ equals to $d_{j}^{i}+...+d_{k}^{i}$) to represent the non-aggregated segment $S^{i}_{j,\,k}$.
@@
FSTA Solution Update. After identifying unstable edges $\{e^{i}_{j_{1}},e^{i}_{j_{2}},...\}$ in each route (which will be addressed in Section 4), where each $e^{i}_{j}$ denotes the edge starting from the $j^{\text{th}}$ node in route $R^{i}$, we break these edges and group the remaining stable edges into segments. To preserve a valid depot, edges connecting to the depot are included in the unstable edge set. After unstable edges are removed, each route $R^{i}$ is then decomposed into multiple disjoint segments $\{x_{0},S^{i}_{1,j_{1}},S^{i}_{j_{1},j_{2}},...\}$, where $x_{0}$ is depot. Each segment ${S}^{i}_{j,\,k}$ is then aggregated into one or two hypernodes $\tilde{S}^{i}_{j,\,k}$, leading to a reduced problem $\tilde{P}$. We then obtain the corresponding solution $\tilde{\mathcal{R}}$ for such reduced problem, where for each $\tilde{R}^{i}\in\tilde{\mathcal{R}}$, we have $\tilde{R}^{i}=(x_{0}\rightarrow\tilde{S}^{i}_{1,j_{1}}\rightarrow\tilde{S}^{i}_{j_{1},j_{2}},...\rightarrow x_{0})$. With fewer nodes than the original problem $P$, re-optimization with a backbone solver becomes more efficient, which is analyzed and confirmed in Appendix A.1. After re-optimization, we obtain a new solution $\tilde{\mathcal{R}}_{+}$ for the reduced problem $\tilde{P}$, which is then recovered into a solution $\mathcal{R}_{+}$ for the original problem $P$ by expanding each hypernode(s) back into its original segment of nodes. This relies on our monotonicity theorem, which guarantees that an improved solution in $\tilde{P}$ maps to an improved solution in $P$.
@@
Theoretical Analysis. We establish a theorem proving FSTA decomposition’s feasibility and monotonicity across multiple VRP variants (e.g. CVRP, CVRPTW), with the proof in Appendix A.2.
Theorem (Feasibility & Monotonicity). If the aggregated solution $\tilde{\mathcal{R}}_{+}$ is feasible to the aggregated problem, then $\mathcal{R}_{+}$ is also feasible to the original, non-aggregated problem.
@@
Moreover, if two feasible aggregated solutions ~+$\tilde{\mathcal{R}}^{1}_{+}$ and ~+$\tilde{\mathcal{R}}^{2}_{+}$ satisfy $f(\tilde{\mathcal{R}}^{1}_{+})\leq f(\tilde{\mathcal{R}}^{2}_{+})$, where $f(\cdot)$ denotes the objective function (total travel cost), their corresponding original solutions also preserve this order: $f(\mathcal{R}^{1}_{+})\leq f(\mathcal{R}^{2}_{+})$.

### v2

@@ Abstract
Iterative heuristics are widely recognized as state-of-the-art for Vehicle Routing Problems (VRPs). In this work, we exploit a critical observation: a large portion of the solution remains stable, i.e., unchanged across search iterations, causing redundant computations, especially for large-scale VRPs with long subtours. To address this, we pioneer the formal study of the First-Segment-Then-Aggregate (FSTA) decomposition technique to accelerate iterative solvers. FSTA preserves stable solution segments during the search, aggregates nodes within each segment into fixed hypernodes, and focuses the search only on unstable portions. Yet, a key challenge lies in identifying which segments should be aggregated. To this end, we introduce Learning-to-Segment (L2Seg), a novel neural framework to intelligently differentiate potentially stable and unstable portions for FSTA decomposition. We present three L2Seg variants: non-autoregressive (globally comprehensive but locally indiscriminate), autoregressive (locally refined but globally deficient), and their synergy. Empirical results on CVRP and VRPTW show that L2Seg accelerates state-of-the-art solvers by 2x to 7x. We further provide in-depth analysis showing why synergy achieves the best performance. Notably, L2Seg is compatible with traditional, learning-based, and hybrid solvers, while supporting various VRPs.
@@ 1 Introduction
Vehicle Routing Problems (VRPs) have profound applications such as in logistics and ride-hailing, driving advances in combinatorial optimization (Laporte, 2009). As NP-hard problems, they are typically tackled with heuristics approximately. Neural Combinatorial Optimization (NCO) (Kool et al., 2018, Bengio et al., 2021, Luo et al., 2024, Berto et al., 2023) has recently introduced machine learning into VRP solving, enabling data-driven decision-making with minimal domain knowledge while matching and even surpassing the performance of meticulously designed heuristics such as Lin-Kernighan-Helsgaun (LKH)(Helsgaun, 2017) and Hybrid Genetic Search (HGS)(Vidal, 2022).
Generally, state-of-the-art VRP solvers predominantly rely on iterative search to refine solutions through local search (e.g., ruin and repair). However, as noted in Section 3, a significant portion of edges stabilize [^1] Specifically, we refer stable edges as those that consistently remain in the solution across iterations, while unstable edges are likely to be re-optimized (see Appendix A.1 for the formal definitions)., or their presence in the solution stops changing between iterations, as the search progresses, despite repeated local search. For example, inner edges of neighboring subtours may remain fixed while only boundary edges undergo frequent combinatorial changes. Intuitively, such stability can be inferred from customer spatial distribution and the solution properties through end-to-end learning. Yet, existing solvers overlook such opportunities, leading to redundant computations that hinder their scalability and efficiency, especially in large-scale VRPs with long subtours.
@@
To identify unstable portions for FSTA decomposition, we then introduce Learning-to-Segment (L2Seg), a novel learning-guided framework that leverages deep models to intelligently differentiate potentially stable and unstable portions, allowing dynamic decomposition for accelerated local search. Realizing this, however, is nontrivial: it involves a large combinatorial decision space requiring accurate segment grouping, and demands modeling complex interdependencies among predicted edges, constraints, spatial distribution, solution structures, and both node and edge features.
@@
To address these challenges, L2Seg proposes encoder-decoder-styled neural models. The encoder integrates graph-level and route-level features using attention and graph neural networks, generating node embeddings that guide edge re-optimization predictions. L2Seg offers three decoders: (1) L2Seg-NAR (Non-Autoregressive): which features one-shot fast global prediction; (2) L2Seg-AR (Autoregressive), which enjoys sequential dependency modeling for high-precision local predictions; and (3) L2Seg-SYN (Synergized), which balances the strengths of both NAR and AR. Notably, this represents a pioneering work that explores the joint decision-making between AR and NAR models in neural combinatorial optimization. Our L2Seg models are trained via a weighted cross-entropy loss on datasets labeled using a lookahead procedure: edge stability is classified based on whether its presence in the solution was changed during iterative re-optimization.
@@
Extensive experiments on large-scale CVRPs and VRPTWs show that L2Seg accelerates backbone heuristics by 2x to 7x, enabling them to outperform state-of-the-art classic, neural, and hybrid baselines, while generalizing well across different customer distributions and problem sizes. Notably, L2Seg exhibits strong flexibility in enhancing various solvers, including the classic LKH-3 Helsgaun (2017) solver, other orthogonal Large Neighborhood Search (LNS) methods Shaw (1998), and learning-guided decomposition method Learning-to-Delegate (L2D) Li et al. (2021). We further analyze the synergy between AR and NAR models, showing their combination achieves the best performance by integrating NAR’s global comprehension with AR’s local precision.
@@
Our contributions are: (1) We make a critical yet underexplored insight that stable segments persist across search iterations in large-scale VRPs, causing redundant computations; (2) We formally study and theoretically prove the properties and applicabilities of First-Segment-Then-Aggregate (FSTA) for various VRPs; (3) We develop Learning-to-Segment (L2Seg), a learning-guided framework with bespoke network architecture, training, and inference for segment identification; (4) We propose autoregressive, non-autoregressive, and their synergistic deep models, pioneering the first-of-its-kind study in NCO; (5) L2Seg consistently accelerates state-of-the-art iterative VRP solvers by 2x to 7x, boosting both classic and learning-based solvers, including other decomposition frameworks.
@@ 2 Related Works
VRP Solvers. Classical VRP solvers include exact methods with guarantees (Baldacci et al., 2012) and practical heuristics (Helsgaun, 2017). Recently, machine learning has been applied to combinatorial optimization, either end-to-end (Kwon et al., 2020, Kool et al., 2018, Fang et al., 2024, Geisler et al., 2022, Gao et al., 2024, Drakulic et al., 2023, Wang et al., 2024, Min et al., 2023, Li et al., 2023a) or learning-guided to unite data-driven insights into human solvers (Li et al., 2021, Lu et al., 2023, Huang et al., 2024; 2023, Hottung et al., 2025). For VRPs, the former could yield competitive performance to classic methods (Luo et al., 2023, Drakulic et al., 2023), while the latter often achieve state-of-the-art performance (Zheng et al., 2024). Among these, most effective VRP solvers rely on iterative search, including classic heuristics such as HGS (Vidal, 2022), LNS (Shaw, 1998) and LKH (Helsgaun, 2017); neural solvers that learn local search (Ma et al., 2021; 2023, Kim et al., 2023, Hottung and Tierney, 2022, Ma et al., 2022); neural constructive solvers integrated with search components (Hottung et al., 2022, Luo et al., 2023, Kim et al., 2021, Sun and Yang, 2023, Chalumeau et al., 2023, Kim et al., 2024, Qiu et al., 2022); and hybrid learning-guided methods like L2D (Li et al., 2021). However, both handcrafted and neural iterative search solvers overlook the redundant computations identified in this paper, particularly in large-scale VRPs.
@@
Decomposition for Large-scale VRPs. Scalability in VRP solvers often relies on effective decomposition that operates on solutions partially (Santini et al., 2023). This includes hand-crafted heuristics, such as LNS (Shaw, 1998) and evolutionary algorithms (Helsgaun, 2017), as well as learning-based methods such as sub-tour grouping (Zong et al., 2022), problem variant reduction (Hou et al., 2023), action space decomposition (Drakulic et al., 2023, Luo et al., 2023, Zhou et al., 2025a) and spatial-based decomposition (Zheng et al., 2024, Zhou et al., 2025b, Pan et al., 2025). In this paper, we present FSTA and L2seg, a fresh [^2] A detailed comparison with representative decomposition methods is provided in Appendix C.1. learning-based decomposition framework that automatically detects unstable edges and aggregates stable segments. Notably, L2Seg holds potential to enhance other decomposition methods, such as LNS (Shaw, 1998) and L2D (Li et al., 2021). While another related work  (Morabit et al., 2024) explores segment stability for re-optimization in a specific dynamic CVRP setting, our work addresses a different problem, i.e., identifying stable segments across search steps to accelerate iterative solvers. And we formally analyze the solution equivalence of FSTA across broader VRP variants. Moreover, L2Seg uniquely designs and integrates three novel deep learning models (AR, NAR, and synergized) to guide FSTA decomposition during search.
@@
AR and NAR Models. In NCO, NAR models make global predictions like edge heatmaps (Sun and Yang, 2023, Li et al., 2023b). However, they struggle to model complex interdependencies, particularly VRP constraints. In contrast, AR models make sequential predictions, e.g., node by node selection in construction solvers (e.g., Luo et al. (2023)). AR offers stronger modeling capacity but might overlook global structure. Recent NCO works combine AR and NAR models in divide-and-conquer frameworks, with NAR for problem splitting and AR for solving (Zheng et al., 2024, Ye et al., 2024, Hou et al., 2023). We are the first to leverage their complementary strengths for joint decision-making, enabling more effective identification of unstable segments in FSTA decomposition.
@@ 3 First-Segment-Then-Aggregate (FSTA)
Figure 1: Percentage of re-optimized edges during iterative search using LKH-3 on 100 CVRP instances. Most edges remain unchanged, suggesting redundant calculations.
@@ 3.1 Vehicle Routing Problems (VRPs)
VRPs aim to minimize total travel costs (often distance or travel time) while serving a set of customers under constraints. Formally, A VRP instance $P$ is defined on a graph $G=\left(V,E\right)$, where each node $x_{i}\in V$ represents a customer and each edge $e_{i,j}\in E$ represents traveling from $x_{i}$ to $x_{j}$ and is associated with a travel cost. For Capacitated VRP (CVRP), vehicles of capacity $C$ start and end at a depot node $x_{0}$. The sum of the demands $d_{i}$ on any route must not exceed $C$, and each customer should be served exactly once. For VRP with Time Windows (VRPTW), each customer is additionally associated with a service time $s_{i}$ and a time window $\bigl[t^{l}_{i},\,t^{r}_{i}\bigr]$ within which service must begin. See Appendix A for the formal definitions of CVRP and VRPTW.
@@ 3.2 FSTA Decomposition
Figure 1 depicts that iterative search solvers perform redundant searches, reoptimizing only a small portion while many edges remain unchanged, especially in large subtours with high capacity $C$. Inspired by Morabit et al. (2024), we formally study the decomposition technique, First-Segment-Then-Aggregate (FSTA), for accelerating iterative search solvers. As shown in the top of Figure 2, FSTA segments the VRP solutions by identifying unstable portions, and then groups them into hypernodes with aggregated attributes. We thus expect more efficient re-optimization on the reduced problems with smaller size. More visualization of FSTA is provided in Appendix B.1.
@@
Segment Definition. Denote the solution (set of routes) of a CVRP as $\mathcal{R}=\{R^{1},R^{2},$...\}, and each route as $R^{i}=(x_{0}\rightarrow x^{i}_{1}\rightarrow x^{i}_{2}\rightarrow...\rightarrow x_{0})\in\mathcal{R}$, where the first and the last nodes in $R^{i}$ are the depot. A segment consists of some consecutive nodes within a route. We denote the segment containing the $j^{\text{th}}$ to $k^{\text{th}}$ nodes of route $i$ as $S^{i}_{j,\,k}=(x_{j}^{i}\rightarrow...\rightarrow x_{k}^{i})$. An aggregated segment $\tilde{S}^{i}_{j,\,k}$ uses one hypernode ($\tilde{S}^{i}_{j,\,k}=\{\tilde{x}^{i}_{j,k}\}$) or two hypernodes ($\tilde{S}^{i}_{j,\,k}=\{\tilde{x}^{i}_{j},\tilde{x}^{i}_{k}\}$) with aggregated attributes (e.g. the demand of $\tilde{x}^{i}_{j,k}$ equals to $d_{j}^{i}+...+d_{k}^{i}$) to represent the non-aggregated segment $S^{i}_{j,\,k}$.
@@
FSTA Solution Update. After identifying unstable edges $\{e^{i}_{j_{1}},e^{i}_{j_{2}},$...\} in each route (which will be addressed in Section 4), where each $e^{i}_{j}$ denotes the edge starting from the $j^{\text{th}}$ node in route $R^{i}$, we break these edges and group the remaining stable edges into segments. To preserve a valid depot, edges connecting to the depot are included in the unstable edge set. After unstable edges are removed, each route $R^{i}$ is then decomposed into multiple disjoint segments $\{x_{0},S^{i}_{1,j_{1}},S^{i}_{j_{1},j_{2}},$...\}, where $x_{0}$ is depot. Each segment ${S}^{i}_{j,\,k}$ is then aggregated into one or two hypernodes $\tilde{S}^{i}_{j,\,k}$, leading to a reduced problem $\tilde{P}$. We then obtain the corresponding solution $\tilde{\mathcal{R}}$ for such reduced problem, where for each $\tilde{R}^{i}\in\tilde{\mathcal{R}}$, we have $\tilde{R}^{i}=(x_{0}\rightarrow\tilde{S}^{i}_{1,j_{1}}\rightarrow\tilde{S}^{i}_{j_{1},j_{2}}...\rightarrow x_{0})$. With fewer nodes than the original problem $P$, re-optimization with a backbone solver becomes more efficient, which is analyzed and confirmed in Appendix B.1. After re-optimization, we obtain a new solution $\tilde{\mathcal{R}}_{+}$ for the reduced problem $\tilde{P}$, which is then recovered into a solution $\mathcal{R}_{+}$ for the original problem $P$ by expanding each hypernode(s) back into its original segment of nodes. This relies on our monotonicity theorem, which guarantees that an improved solution in $\tilde{P}$ maps to an improved solution in $P$.
@@
Theoretical Analysis. We establish a theorem proving FSTA’s feasibility and monotonicity across multiple VRP variants (e.g. CVRP, VRPTW, VRPB, and 1-VRPPD), with the proof in Appendix B.2.
@@
Theorem (Feasibility & Monotonicity). If the aggregated solution $\tilde{\mathcal{R}}_{+}$ is feasible to the aggregated problem, then $\mathcal{R}_{+}$ is also feasible to the original, non-aggregated problem. Moreover, if two feasible aggregated solutions $\tilde{\mathcal{R}}^{1}_{+}$ and $\tilde{\mathcal{R}}^{2}_{+}$ satisfy $f(\tilde{\mathcal{R}}^{1}_{+})\leq f(\tilde{\mathcal{R}}^{2}_{+})$, where $f(\cdot)$ denotes the objective function (total travel cost), their corresponding original solutions also preserve this order: $f(\mathcal{R}^{1}_{+})\leq f(\mathcal{R}^{2}_{+})$.